# Bird’s-Eye View to Street-View: A Survey

### v1

@@ Abstract
Index Terms:
Computer Society, IEEE, IEEEtran, journal, , paper, template.
I Introduction
@@
Generating street-view images from satellite-view is one of the rising topics in the cross-view image translation domain. It aims to generate geometrically consistent street views corresponding to the center-located area on the provided satellite view. As satellite images are becoming more widely available (e.g., Google Earth), it is becoming easier to cover almost every spot of the globe. However, such a capability does not exist for street-view images [1], as the majority of street-views are captured around known landmarks and major roads [2][3]. It helps enhance the content in areas that are expensive or difficult to reach by people or vehicles. Also, street-view images enable more accurate geo-localization than satellite-view images.
@@
Street-view synthesis is helpful for various downstream tasks, such as image geo-localization, which is considered a fundamental problem in many applications such as robotics, augmented reality, and autonomous driving [4]. The retrieval component is limited to identifying and locating similar objects, so adopting satellite to street-view synthesis plays a vital role in this task, where the generator can learn more valuable features that help in precise location retrieval. Moreover, street-view images have many additional useful details not in satellite images, like building facades [4]. Furthermore, cross-view matching between satellite and street views aids in predicting street-view segmentation maps based on satellite views of the same location for geo-orientation and geo-location estimation tasks by extracting features from a satellite image and mapping them into street-view, conveying semantics from street-view to satellite view [5]. Moreover, the representation derived from the conditional Generative Adversarial Networks (GANs) for generating street views from satellite images aids in distinguishing between rural and urban land covers. [6].
@@
In recent years, synthesizing street-view from satellite images has been investigated by Generative Adversarial Network-based methods [7][8][6][9][4], or by combining U-Net [10] and BicycleGAN such as [1]. Differently, [11] and [8] explored the conditional GAN that learns the generation of both images, either street-view or satellite-view, and their corresponding semantic maps simultaneously. In addition, [12] and [13] investigated the semantic-guided cGAN where the semantic map concatenated with the input image. As a result, the semantic map serves as a guide to direct the model on generating images in another domain. On the other hand, street-view to satellite synthesis, has been explored by employing conditional GAN [14] for street-view geo-localization task for predicting street-view location. In addition, some other works investigated the street-view geo-localization task by matching it with satellite images using Convolutional Neural Networks (CNN) and feature matching-based methods [15][2], Siamese Network [16][17][18], and Siamese and Triplet Network [19].
@@
Figure 1 illustrates a general overview of the satellite to street-view generation. As a result of the challenging nature of this problem, it leaves an opportunity for further improvements. Despite the progress of satellite-to-street-view synthesis, it has limitations, as a satellite image only shows the top surface of an object. Therefore, it led to a hallucinatory representation of the facades of the objects. In addition, indistinguishable small objects such as cars, tree shadows, sidewalks, street lines, and fences are still being ignored in street-view synthesis.
@@
II Bird Eye View to Street View Methods
II-A Conditional GAN-based Models
@@
II-A1 X-Fork [11]
: Regmi and Borji introduced two architectures based on conditional GANs called Cross-view Fork (X-Fork) and Sequential-view (Seq-X) for cross-image synthesis from street view to aerial view and vice versa. The X-Fork architecture used the baseline discriminator, where the generator network is forked, to produce segmentation maps and images, as shown in Figure 3(a). The Seq-X architecture uses two conditional GAN networks: the first network generates baseline cross-view images, which are fed to the second network as conditioning input to create the segmentation map in the same view, as illustrated in Figure 3(b). The results illustrated that X-Fork achieved higher accuracy with Top-1 and Top-5 for 64$\times$64 resolution. On the other hand, X-Seq performed better with 256$\times$256 resolution. This study employs two datasets: the first is the Dayton [19], and the second is the CVUSA.
II-A2 BEVGen [20]
: It proposed a conditional GAN-based model for generating realistic street-view images from Bird’s Eye View layout for traffic scenarios. The evaluation is conducted on the NuScenes 360 driving dataset [21]. Figure 3(c) shows an overview of BEVGen architecture, where it encoded both street-view images and BEV layout to latent discrete representation and flattened them for the autoregressive transformer. Subsequently, the spatial embeddings are added to the street view and Birds Eye View (BEV) tokens, and the pairwise bias is combined with the attention weight. After that, the decoder receives the tokens to get the generated street-view images. BEVGen compares with the baseline model, which has the same GPT architecture and uses identical first-stage encoders and decoders. The result indicates that BEVGen outperformed the baseline on metrics such as Fréchet Inception Distance (FID), road mIoU, and vehicle mIoU. As a limitation, BEVGen [20] faces challenges in accurately generating small objects like people and specific vehicles.
@@ II-A2 BEVGen [20]
(c) BEVGen [20] (d) GeoUrban [1]
Refer to caption Refer to caption
(e) FeatureFusion [14] (f) cGANs [6]
@@ II-A3 GeoUrban* [1]:
It exploits geometry to synthesize street-view images from satellite images. This study shows a 5km by 5km region of interest in the center of London City. The ground truth of depth and semantic satellite images is generated from stereo matching and supervised classification with post-correction. The architecture consists of three stages, as shown in Figure 3(d). The first is the satellite stage, which uses U-Net [10] to calculate a given satellite image’s depth and semantic images. The second stage is the geo-transformation stage, which converts the semantic and depth images of satellite images into their corresponding semantic and depth panorama images. The third stage, the street-view stage, is another U-Net [10] that refines the semantic panorama. Eventually, the BicycleGAN [22] is applied to generate realistic images from the semantic panorama. The novelty of the proposed method lies in offering a new transformation module for generating street-view images constrained by the geometric information of satellite images. Also, a novel weighted mask is utilized to mitigate the differences between the geometric panorama and satellite-depth images. The results demonstrated that the GeoUrban [1] outperformed other methods like Pix2Pix [23] and Regmi et al. [11]’s model with an accuracy of 72.9%.
II-A4 FeatureFusion [14]:
@@
It takes advantage of conditional GANs to synthesize aerial images from ground-level panoramas. The X-Seq [11] is adopted with some modifications, such as reshaping the features into $2\times 2$ and then performing different up-convolution operations to generate $512\times 512$ aerial image resolution. The ground-level image and its edge-map are inputs to the X-Fork generator [11], as shown in Figure 3(e), which generates a synthesized aerial image and its corresponding edge-map. After that, FeatureFusion learns the joint feature representation of the ground-level image and the synthesized aerial image, and the features of the three images are integrated, followed by a fully connected layout to get a robust representation that matches the aerial feature representation. The novelty of the method lies in combining the edge map and semantic segmentation with input images to enhance the synthesis of cross-view. The results showed that the proposed feature fusion-based outperformed other approaches, including [2, 5, 19, 17] with 48.75% top-1 recall, 81.27% top-10 recall, and 95.98% top-1 accuracy. Moreover, FeatureFusion [14] experiments on the CVUSA and a new dataset the authors collected, the Orlando-Pittsburgh (OP) dataset.
II-A5 cGANs [6]
@@
use conditional GANs for generating ground-level images from satellite images and investigated cGANs feature maps for the land-cover classification. The satellite photos are collected via the Google Map Statistics API, while the ground-level images are taken from the Geograph API with known locations. The proposed framework addresses the problem of expressing an aerial picture’s information on how things appear on a ground-level image so that it can be used as input to the cGANs to generate ground-level photos. However, feeding 2D satellite image vectors to cGAN is ineffective due to the lack of structural similarity between satellite and ground-level images. Therefore, three types of image patch embedding are utilized to capture structural similarity: grayscale, HSV, and CNN. Figure 3(f) shows the proposed framework where the 1D image embedding vector is joined with a noise vector for the generator’s input. The feature maps generated by cGANs achieved the highest accuracy, with a value of 73.14%.
II-B Multi Generator and Discriminator Models
@@
II-B1 BridgeGAN [24]:
A multi-GAN model for generating bird-view images, BridgeGAN, introduces an intermediate view called the homography view to narrow the gap between the bird view and the frontal view. BridgeGAN consists of three GAN networks, as shown in Figure 4(a), each representing a specific view domain, namely homography-view, bird-view, and frontal-view. Each view domain has three components: a generator, a discriminator, and an encoder. In addition, shared layers are used to learn integrated intermediate representations since all three domains have the same semantic meaning. Cycle consistency loss consists of backward and forward cycles, while the introduced RestNet50 loss network imposes cross-view feature consistency. The novelty of BridgeGAN lies in addressing bird-view image generation from frontal-view images. The BridgeGAN achieves the best performance with values of 0.596 for SSIM, 5.01 for PSNR, and 0.242 for LPIPS compared with other baselines: Pix2Pix [23], cycleGAN [25], DiscoGAN [26], and CoGAN [27]. The dataset is obtained by collecting images from the Grand Theft Auto V video game using Palazzi et al.’s [28] framework.
II-B2 SelectionGAN [13]:
A multi-channel attention selection GAN comprises two stages for synthesizing images conditioned on the reference image and semantic map, as illustrated in Figure 4(b). The first stage focuses on the semantic features of the views, takes the target semantic map and the condition image (i.e., the ground-level image), and produces initial coarse results. The second stage targets the appearance details by refining the initial results via the multi-channel attention selection mechanism. The novelty lies in using coarse-to-fine inference to investigate cascaded semantic guidance. Moreover, a multi-attention module is offered to choose the intermediate generation attentively, which helps improve output quality. SelectionGAN [13] model is evaluated on three datasets: CVUSA [2] for aerial to ground-level images, Dayton [19] for ground-level to aerial images, and Ego2Top [29] (consists of pairs of two images of the same scene from two different views). The SelectionGAN achieved the best performance compared to pix2pix [23], X-Fork [11], X-Seq [11], and PGSL [5]. In some cases, on the CVUSA dataset, the X-Seq [11] occasionally performs slightly higher; however, SelectionGAN generates more realistic images.
@@ II-B2 SelectionGAN [13]:
Refer to caption Refer to caption
(c) PAGAN [8] (d) LGGAN [12]
Figure 4: Multi GAN-based models.
II-B3 PAGAN [8]:
Progressive Attention GAN (PAGAN) has two primary components: a cross-stage attention module and a multi-stage progressive generation framework. Figure 4(c) shows an overview of the PAGAN, in which cross-stage attention is applied to generate a panorama through several coherent stages. PAGAN [8] comprises multiple segmentation and image generators for generating the street-view panorama and its segmentation map. In the first stage, the aerial image is fed to the image generator and segmentation generator, which produce the panorama and its corresponding segmentation map. After that, the generated image is refined by applying the cross-stage attention module, which forwards the semantic information stage by stage. The novelty of PAGAN lies in proposing a progressive generative framework for the high-resolution generation of street-view panoramas. Also, introducing a new cross-stage attention mechanism to bridge the subsequent generational stages allows the quality of street-view panoramas to improve continuously. The PAGAN outperformed other recent works: Pix2Pix [23], X-Fork [11], X-Seq [11], SelectionGAN [13], GAS2G [1], especially in terms of KL and FID scores, where the PAGAN reduced these scores significantly. In this study, two datasets are used, namely, the CVUSA and Orlando-Pittsburgh (OP) datasets.
II-B4 LGGAN [12]:
A Local class-specific and Global image-level GANs (LGGAN) model addresses semantic-guided-based scene generation tasks. LGGAN generation module is composed of three parts: a class-specific local generator for enriching the local scene details, an image-level global generator for capturing target images’ overall structure and layout, and a weight-map generator for combining the resulted images from the global and local generators with the learned weight maps, as shown in Figure 4(d). The discriminator tries to differentiate generated images from semantic and image space. In addition, a new classification module is proposed to learn better discriminative class-specific feature representation. The novelty of LGGAN is that it considers both local and global contexts. The LGGAN performs better on most evaluation metrics compared with other recent works, such as pix2pix [23], SelectionGAN [13], X-Seq [11], and X-Fork [11]. While SelectionGAN achieved slightly better performance on pixel-level metrics, i.e., PSNR, SD, and SSIM, than LGGAN. However, LGGAN generates more visually detailed objects like trees, buildings, and cars than SelectionGAN. The image translation part has been evaluated using the Dayton and CVUSA datasets.
II-C Image-to-Image Translation Frameworks
@@
II-C1 Pix2Pix [23]
: A general framework that investigates using the conditional GAN for image-to-image translation. Though Pix2Pix [23] is not specific to bird-view to street-view, it has been extensively cited in this domain. The Pix2Pix [23] method is evaluated on many tasks and datasets: i) semantic labels to photos trained on Cityscapes [30], ii) map to aerial employs Google Map images, iii) edges to photo for [31] and [32], iv) sketch to photo for human sketches utilizing [33], and v) day to night using [34] data. The architecture of the generator and discriminator has been adapted from [35], where both are composed of Conv-BN-ReLU [36]. The skip connections are added to the generator, following the general shape of the U-Net [10], to allow low-level information from initial layers. Moreover, a novel discriminator, termed PatchGAN, penalizes structure at the image patch scale. The results showed that Pix2Pix [23] achieved the second-best result on the colorization task compared with Colorful Colorization [37], and a variant of their approach employed $\ell_{2}$ regression [37]. Additionally, the Pix2Pix [23] model’s aerial image generated from the map deceived the participants far more than the baseline. On the other hand, the generated map from an aerial image misled the participants by 6.1%, which is not substantially exceeding the baseline.
II-D Semantic Feature Learning and Representation-based Models

### v2

@@ Abstract
keywords:
street-view synthesis, cross-view translation, bird-eye-view, satellite images.
1 Introduction
@@
Generating street-view images from satellite-view is one of the rising topics in the cross-view image translation domain. It aims to generate geometrically consistent street views corresponding to the center-located area on the provided satellite view. As satellite images are becoming more widely available (e.g., Google Earth), it is becoming easier to cover almost every spot of the globe. However, such a capability does not exist for street-view images [32], as the majority of street views are captured around known landmarks and major roads [53][10]. It helps enhance the content in areas that are expensive or difficult to reach by people or vehicles. Also, street-view images enable more accurate geo-localization than satellite-view images.
@@
Street-view synthesis is helpful for various downstream tasks, such as image geo-localization, which is considered a fundamental problem in many applications such as robotics, augmented reality, and autonomous driving [48]. The retrieval component is limited to identifying and locating similar objects, so adopting satellite to street-view synthesis plays a vital role in this task, where the generator can learn more valuable features that help in precise location retrieval. Moreover, street-view images have many additional useful details not in satellite images, such as building facades [48]. Furthermore, cross-view matching between satellite and street views helps in predicting street-view segmentation maps based on satellite views of the same location for geo-orientation and geo-location estimation tasks by extracting features from a satellite image and mapping them into street-view, conveying semantics from street-view to satellite view [57]. Moreover, the representation derived from the conditional Generative Adversarial Networks (GANs) for generating street views from satellite images aids in distinguishing between rural and urban land covers [12].
@@
In recent years, synthesizing street-view from satellite images has been investigated by Generative Adversarial Network-based methods [42][54][12][38][48], or by combining U-Net [39] and BicycleGAN, such as [32]. Differently, [36] and [54] explored the conditional GAN that learns the generation of both images, either street-view or satellite-view, and their corresponding semantic maps simultaneously. In addition, [47] and [46] investigated the semantic-guided cGAN where the semantic map is concatenated with the input image. As a result, the semantic map serves as a guide to direct the model in generating images in another domain. On the other hand, street-view to satellite synthesis, has been explored by employing a conditional GAN [37] for street-view geo-localization task for predicting street-view location. In addition, some other works investigated the street-view geo-localization task by matching it with satellite images using Convolutional Neural Networks (CNN) and feature matching-based methods [43][53], Siamese Network [29][16][30], and Siamese and Triplet Network [50]. Recently, more advanced methods are being used for satellite to street-view synthesis domain, such as transformer-based [45] and diffusion-based methods [28], [26].
@@
Figure 1 illustrates a general overview of the satellite to street-view generation. As a result of the challenging nature of this problem, it leaves an opportunity for further improvements. Despite the progress of satellite-to-street-view synthesis, it has limitations, as a satellite image only shows the top surface of an object. Therefore, it led to a hallucinatory representation of the facades of the objects. In addition, indistinguishable small objects such as cars, tree shadows, sidewalks, and fences are still being ignored in street-view synthesis.
@@
Figure 3: Taxonomy diagram of bird eye view to street-view existing methods, showing the classification and corresponding methods within each category.
2 Bird Eye View to Street View Methods
This section reviews various methods in the bird-eye view to street-view domain. We categorize each method based on the applied techniques. Figure 3 provides an overview for these methods and their corresponding categories, and Table 1 presents a summary of the reviewed methods.
2.1 Conditional GAN-based Models
@@
X-Fork [36]: Regmi and Borji introduced two architectures based on conditional GANs called Cross-view Fork (X-Fork) and Sequential-view (Seq-X) for cross-image synthesis from street view to aerial view and vice versa. The X-Fork architecture used the baseline discriminator, where the generator network is forked, to produce segmentation maps and images, as shown in Figure 4(b). The Seq-X architecture uses two conditional GAN networks: the first network generates baseline cross-view images, which are fed to the second network as conditioning input to create the segmentation map in the same view, as illustrated in Figure 4(a). The results illustrated that X-Fork achieved higher accuracy with Top-1 and Top-5 for $\times64$ resolution. On the other hand, X-Seq performed better with $\times256$ resolution. This study employs two datasets: the first is the Dayton [50], and the second is the CVUSA.
@@
(d) FeatureFusion [37] (e) cGANs [12]
(f) PanoGAN [55]
@@
GeoUrban* [32]: It exploits geometry to synthesize street-view images from satellite images. This study shows a 5km by 5km region of interest in the center of London City. The ground truth of depth and semantic satellite images is generated from stereo matching and supervised classification with post-correction. The architecture consists of three stages, as shown in Figure 4(c). The first is the satellite stage, which uses U-Net [39] to calculate a given satellite image’s depth and semantic images. The second stage is the geo-transformation stage, which converts the semantic and depth images of satellite images into their corresponding semantic and depth panorama images. The third stage, the street-view stage, is another U-Net [39] that refines the semantic panorama. Eventually, the BicycleGAN [64] is applied to generate realistic images from the semantic panorama. The novelty of the proposed method lies in offering a new transformation module for generating street-view images constrained by the geometric information of satellite images. Also, a novel weighted mask is utilized to mitigate the differences between the geometric panorama and satellite-depth images. The results demonstrated that the GeoUrban [32] outperformed other methods like Pix2Pix [18] and Regmi et al. [36]’s model with an accuracy of 72.9%.
@@
FeatureFusion [37]: It takes advantage of conditional GANs to synthesize aerial images from ground-level panoramas. The X-Seq [36] is adopted with some modifications, such as reshaping the features into $2\times 2$ and then performing different up-convolution operations to generate $512\times 512$ aerial image resolution. The ground-level image and its edge-map are inputs to the X-Fork generator [36], as shown in Figure 4(d), which generates a synthesized aerial image and its corresponding edge-map. After that, FeatureFusion learns the joint feature representation of the ground-level image and the synthesized aerial image, and the features of the three images are integrated, followed by a fully connected layout to get a robust representation that matches the aerial feature representation. The novelty of the method lies in combining the edge map and semantic segmentation with input images to enhance the synthesis of cross-view. The results showed that the proposed feature fusion-based outperformed other approaches, including [53, 57, 50, 16] with 48.75% top-1 recall, 81.27% top-10 recall, and 95.98% top-1 accuracy. Moreover, FeatureFusion [37] experiments on the CVUSA and a new dataset the authors collected, the Orlando-Pittsburgh (OP) dataset.
@@
cGANs [12]: use conditional GANs for generating ground-level images from satellite images and investigated cGANs feature maps for the land-cover classification. The satellite photos are collected via the Google Map Statistics API, while the ground-level images are taken from the Geograph API with known locations. The proposed framework addresses the problem of expressing an aerial picture’s information on how things appear on a ground-level image so that it can be used as input to the cGANs to generate ground-level photos. However, feeding 2D satellite image vectors to cGAN is ineffective due to the lack of structural similarity between satellite and ground-level images. Therefore, three types of image patch embedding are utilized to capture structural similarity: grayscale, HSV, and CNN. Figure 4(e) shows the proposed framework where the 1D image embedding vector is joined with a noise vector for the generator’s input. The feature maps generated by cGANs achieved the highest accuracy, with a value of 73.14%.
PanoGAN [54]:Wu et al. [54] proposed an innovative adversarial feedback GAN-based framework (PanoGAN) designed satellite to street-view synthesis task. It consists of two main components: a dual-branch discriminator and adversarial feedback module. During the training phase, the segmentation maps have been incorporated into the proposed framework to guide the street-view generation process and enhance the geometric consistency. PanoGAN has been designed in a way to improve the generator-discriminator cooperation during the generation process. First, the satellite image is input into the generator to generate the panorama street-view image and its corresponding segmentation map. Then, the discriminator’s feature responses are encoded through the adversarial model and fed it back into the generator, to produce the refined version of panorama street-view and segmentation map image pair. This iterative process allows continuous improvements by leveraging discrimination feedback as a powerful source for enhancing the panorama street-view images and their associated segmentation maps. The novelty of this work lies in proposing the dual-branch discrimination strategy to ensure semantic consistency and high-quality images. In addition, the adversarial feedback module is introduced to improve the generation process. The results illustrate that, PanoGAN reached to the highest scores on CVUSA dataset with SSIM, PSNR, and KL values of 0.4437, 20.9467, and 4.20±1.19, respectively. On Orlando-Pittsburgh (OP) dataset it reached to the highest scores on OP dataset in terms of SSIM, KL and SD with values of 0.4428 , 7.89±1.12, and 18.0543, respectively. Figure 4 (f) shows an overview of PanoGAN architecture.
2.2 Multi Generator and Discriminator Models
@@
BridgeGAN [65]: A multi-GAN model for generating bird-view images, BridgeGAN, introduces an intermediate view called the homography view to narrow the gap between the bird view and the frontal view. BridgeGAN consists of three GAN networks, as shown in Figure 5(a), each representing a specific view domain, namely homography-view, bird-view, and frontal-view. Each view domain has three components: a generator, a discriminator, and an encoder. In addition, shared layers are used to learn integrated intermediate representations since all three domains have the same semantic meaning. Cycle consistency loss consists of backward and forward cycles, while the introduced RestNet50 loss network imposes cross-view feature consistency. The novelty of BridgeGAN lies in addressing bird-view image generation from frontal-view images. The BridgeGAN achieves the best performance with values of 0.596 for SSIM, 5.01 for PSNR, and 0.242 for LPIPS compared with other baselines: Pix2Pix [18], CycleGAN [63], DiscoGAN [22], and CoGAN [31]. The dataset is obtained by collecting images from the Grand Theft Auto V video game using Palazzi et al.’s [34] framework.
SelectionGAN [46]: A multi-channel attention selection GAN comprises two stages for synthesizing images conditioned on the reference image and semantic map, as illustrated in Figure 5(b). The first stage focuses on the semantic features of the views, takes the target semantic map and the condition image (i.e., the ground-level image), and produces initial coarse results. The second stage targets the appearance details by refining the initial results via the multi-channel attention selection mechanism. The novelty lies in using coarse-to-fine inference to investigate cascaded semantic guidance. Moreover, a multi-attention module is offered to choose the intermediate generation attentively, which helps improve output quality. SelectionGAN [46] model is evaluated on three datasets: CVUSA [53] for aerial to ground-level images, Dayton [50] for ground-level to aerial images, and Ego2Top [1] (consists of pairs of two images of the same scene from two different views). The SelectionGAN achieved the best performance compared to Pix2Pix [18], X-Fork [36], X-Seq [36], and PGSL [57]. In some cases, on the CVUSA dataset, the X-Seq [36] occasionally performs slightly higher; however, SelectionGAN generates more realistic images.
@@
(c) PAGAN [54]
(d) LGGAN [47]
Figure 5: Multi generator and discriminator models.
PAGAN [54]: Progressive Attention GAN has two primary components: a cross-stage attention module and a multi-stage progressive generation framework. Figure 5(c) shows an overview of the PAGAN, in which cross-stage attention is applied to generate a panorama through several coherent stages. PAGAN [54] comprises multiple segmentation and image generators for generating the street-view panorama and its segmentation map. In the first stage, the aerial image is fed to the image generator and segmentation generator, which produce the panorama and its corresponding segmentation map. After that, the generated image is refined by applying the cross-stage attention module, which forwards the semantic information stage by stage. The novelty of PAGAN lies in proposing a progressive generative framework for the high-resolution generation of street-view panoramas. Also, introducing a new cross-stage attention mechanism to bridge the subsequent generational stages allows the quality of street-view panoramas to improve continuously. The PAGAN outperformed other recent works: Pix2Pix [18], X-Fork [36], X-Seq [36], SelectionGAN [46], GAS2G [32], especially in terms of KL and FID scores, where the PAGAN reduced these scores significantly. In this study, two datasets are used, namely, the CVUSA and Orlando-Pittsburgh (OP) datasets.
LGGAN [47]: A Local class-specific and Global image-level GANs (LGGAN) model addresses semantic-guided-based scene generation tasks. LGGAN generation module is composed of three parts: a class-specific local generator for enriching the local scene details, an image-level global generator for capturing target images’ overall structure and layout, and a weight-map generator for combining the resulted images from the global and local generators with the learned weight maps, as shown in Figure 5(d). The discriminator tries to differentiate generated images from semantic and image space. In addition, a new classification module is proposed to learn better discriminative class-specific feature representation. The novelty of LGGAN is that it considers both local and global contexts. The LGGAN performs better on most evaluation metrics compared with other recent works, such as pix2pix [18], SelectionGAN [46], X-Seq [36], and X-Fork [36]. While SelectionGAN achieved slightly better performance on pixel-level metrics, i.e., PSNR, SD, and SSIM, than LGGAN. However, LGGAN generates more visually detailed objects like trees, buildings, and cars than SelectionGAN. The image translation part has been evaluated using the Dayton and CVUSA datasets.
2.3 Image-to-Image Translation Frameworks
@@
Pix2Pix [18]: A general framework that investigates using the conditional GAN for image-to-image translation. Though Pix2Pix [18] is not specific to bird-view to street-view, it has been extensively cited in this domain. The Pix2Pix [18] method is evaluated on many tasks and datasets: i) semantic labels to photos trained on Cityscapes [9], ii) map to aerial employs Google Map images, iii) edges to photo for [62] and [56], iv) sketch to photo for human sketches utilizing [13], and v) day to night using [24] data. The architecture of the generator and discriminator has been adapted from [35], where both are composed of Conv-BN-ReLU [17]. The skip connections are added to the generator, following the general shape of the U-Net [39], to allow low-level information from initial layers. Moreover, a novel discriminator, termed PatchGAN, penalizes structure at the image patch scale. The results showed that Pix2Pix [18] achieved the second-best result on the colorization task compared with Colorful Colorization [59], and a variant of their approach employed $\ell_{2}$ regression [59]. Additionally, the Pix2Pix [18] model’s aerial image generated from the map deceived the participants far more than the baseline. On the other hand, the generated map from an aerial image misled the participants by 6.1%, which is not substantially exceeding the baseline.
2.4 Representation-based Models